OpenStudy (anonymous):

15-7(sqrt(15)) / 15+7(sqrt(15)) how can I simplify this? I know it comes to: - (32 - 7(sqrt( 15 )) / 17 But can someone show their work? I can figure it out if I can see the steps. Please??? :)

OpenStudy (anonymous):

The "conjugate" of \[15+7\sqrt{15}\] is \[15-7\sqrt{15}\] Multiplying those two together produces a rational value. -- Try doing this: \[\frac{15-7\sqrt{15}}{15+7\sqrt{15}} \cdot \frac{15-7\sqrt{15}}{15-7\sqrt{15}}\] Don't cross out the opposites (or we would be back where we started!. Just multiply them together: \[\frac{(15-7\sqrt{15})(15-7\sqrt{15})}{(15+7\sqrt{15})(15-7\sqrt{15})} \]

OpenStudy (anonymous):

@student1988... @muntoo just gave you that expert help you wanted.

OpenStudy (anonymous):

hey thanks! I did this, but my multiplication was wrong.